# AutoDAN: Generating Stealthy Jailbreak Prompts on Aligned Large Language Models

### v1

@@ 1 Introduction
However, there are two limitations of existing jailbreak methods: Firstly, automatic attacks like GCG  inevitably request a search scheme guided by the gradient information on tokens. Although it provides a way to automatically generate jailbreak prompts, this leads to an intrinsic drawback: they often generate jailbreak prompts composed of nonsensical sequences or gibberish, i.e., without any semantic meaning (Morris et al., 2020). This severe flaw makes them highly susceptible to naive defense mechanisms like perplexity-based detection. As a recent study (Jain et al., 2023) has demonstrated, such straightforward defense can easily identify these nonsensical prompts and completely undermine the attack success rate of the GCG attack. Secondly, despite the fact that manual attacks can discover stealthiness jailbreak prompts, the jailbreak prompts are often handcrafted by individual LLM users, therefore facing scalability and adaptability challenges. Moreover, such methods may not adapt quickly to updated LLMs, reducing their effectiveness over time (Albert, 2023; ONeal, 2023). Hence, a natural question emerges: “Is it possible to automatically generate stealthy jailbreak attacks? ”
@@ 3.1 Preliminaries
3.2 AutoDAN
As discussed in the introduction, although current jailbreak methods have limitations, when viewed from a different perspective, existing handcrafted methods have offered a valuable starting point, while the automatic method has introduced a valuable score function that guides the optimization of prompts. Thus, we utilize handcrafted jailbreak prompts, such as the DAN series, as an initial point for our semantically meaningful jailbreak prompt. This approach significantly benefits the initialization of our method since it enables the exploration of jailbreak prompts in a space that is more proximate to the potential solution. Once we have an initialized searching space, it becomes appropriate to employ evolutionary algorithms (e.g., genetic algorithm) based on score function, to identify jailbreak prompts that can effectively compromise the victim LLM.
Thus, our primary objective is to develop a genetic algorithm that utilizes handcrafted prompts as its initialization point and jailbreak attack loss as its scoring function. However, before implementing such a method, we must address three questions: (1) How do we define an initialization space derived from the provided handcrafted prompt? (2) How to design the fitness score to guide the optimization? (3) More importantly, in a discrete representation space of text, how do we formulate effective policies for sample selection, crossover, and mutation?
@@ 3.3 Population Initialization
Initialization policy plays a pivotal role in genetic algorithms because it can significantly influence the algorithm’s convergence speed and the quality of the final solution. To answer the first question, i.e., design an effective initialization policy for AutoDAN, we should bear in mind two key considerations: 1) The prototype handcrafted jailbreak prompt has already demonstrated efficacy in specific scenarios, making it a valuable foundation; thus, it is imperative not to deviate too far from it. 2) Ensuring the diversity of the initial population is crucial, as it prevents premature convergence to sub-optimal solutions and promotes a broader exploration of the solution space. To preserve the basic features of the prototype handcrafted jailbreak prompt while also promoting diversification, we employ LLMs as the agents responsible for revising the prototype prompt, as illustrated in Alg. 4. The rationale behind this scheme is that the modifications proposed by LLM can preserve the inherent logical flow and meaning of the original sentences, while simultaneously introducing diversity in word selection and sentence structure.
3.4 Fitness evaluation.
To answer the second question, since the goal of jailbreak attacks can be formulated as Eq. 2, we can directly use a function that calculates this likelihood for evaluating the fitness of the individuals in genetic algorithms. Here, we adopt the log-likelihood that was introduced by Zou et al. (2023) as the loss function, namely, given a specific jailbreak prompt $J_{i}$, the loss can be calculated by:
@@
3.4.1 AutoDAN-GA.
Based on the initialization scheme and fitness evaluation function, we can further use the genetic algorithm to conduct the optimization. The core of the genetic algorithm is to design the crossover and mutation functions. By using a multi-point crossover scheme, we can develop our first version genetic algorithm AutoDAN-GA. We provide the detailed implementation of AutoDAN-GA in Appendix C since, here, we hope to discuss how to formulate more effective policies for handling the structural discrete text data, by using its intrinsic characteristics.
3.5 Hierarchical Genetic Algorithm for Structured Discrete Data
@@
Given the population that is initialized by Alg. 4, the proposed AutoDAN will first evaluate the fitness score for every individual in the population following Eq. 3. After the fitness evaluation, the next step is to choose the individuals for crossover and mutation. Let’s assume that we have a population containing $N$ prompts. Given an elitism rate $\alpha$, we first allow the top $N*\alpha$ prompts with the highest fitness scores to directly proceed to the next iteration without any modification, which ensures the fitness score is consistently dropping. Then, to determine the remaining $N-N*\alpha$ prompts needed in the next iteration, we first use a selection method that chooses the prompt based on its score. Specifically, the selection probability for a prompt $J_{i}$ is determined using the softmax function:
@@
After the selection process, we will have $N-N*\alpha$ “parent prompts” ready for crossover and mutation. Then for each of these prompts, we perform a multi-point crossover at a probability $p_{crossover}$ with another parent prompt. The multi-point crossover [^1] Since the multi-point crossover is straightforward, we defer the detailed description to Appendix C. scheme can be summarized as exchanging the sentences of two prompts at multiple breakpoints. Subsequently, the prompts after crossover will be conducted a mutation at a probability $p_{mutation}$. We let the LLM-based diversification introduced in Alg. 4 to also serve as the mutation function due to its ability to preserve the global meaning and introduce diversity. We delineate the aforementioned process in Alg. 6. For the $N-N*\alpha$ selected data, this function returns $N-N*\alpha$ offsprings. Combining these offsprings with the elite samples that we preserve, we will get $N$ prompts in total for the next iteration.
@@
At the sentence level, the search space is primarily around the word choices. After scoring each prompt using the fitness score introduced in Eq. 3, we can assign the fitness score to every word present in the corresponding prompt. Since one word can appear in multiple prompts, we set the average score as the final metric to quantify the significance of each word in achieving successful attacks. To further consider the potential instability of the fitness score in the optimization process, we incorporate a momentum-based design into the word scoring, i.e., deciding the final fitness score of a word based on the average number of the score in current iteration and the last iteration. As detailed in Alg. 7, after filtering out some common words and proper nouns (line 4) , we can obtain a word score dictionary (line 6). From this dictionary, we choose the words with top $K$ scores to replace their near-synonym in other prompts, as described in Alg. 8.
Termination criteria. To ensure the effectiveness and efficiency of AutoDAN, we adopt termination criteria combined with a max iteration test and refusal signals test. If the algorithm has exhausted the maximum number of iterations or no keyword in a set list $L_{refuse}$ is detected in the LLM’s top $K$ words of its response, AutoDAN will terminate and return the current optimal jailbreak prompt, which has the highest fitness score. Finally, the AutoDAN-HGA can be delineated in Alg. 1.
@@
1:Input Prototype jailbreak prompt $J_{p}$, hyper-parameters
2:Initialize population with LLM-based Diversification in Alg. 4
3:Evaluate the fitness score of each individual in population following Eq. 3
4:for $iteration$ in paragraph-level Iterations do
@@
6:     if model responses have no word in $L_{refuse}$ then
7:         return Optimal jailbreak prompt $J_{max}$ with highest fitness score
8:     Select elite and parent prompts following Eq.4
9:     Conduct crossover and mutation on parent prompts following Alg. 6
10:     for $iteration$ in sentence-level Iterations do
11:         Evaluate the fitness score of each individual in population following Eq. 3
12:         if model responses have no word in $L_{refuse}$ then Break
13:         Calculate momentum word score by Alg. 7
14:         Update sentences in each prompt by Alg. 8
15:     Evaluate the fitness score of each individual in population following Eq. 3
16:     if model responses have no word in $L_{refuse}$ then
17:         return Optimal jailbreak prompt $J_{max}$ with highest fitness score
18:return Optimal jailbreak prompt $J_{max}$ with highest fitness score
@@ 4.1 Experimental Setups
Metrics. We use two metrics to evaluate the performance of jailbreak methods. The first metric is the keyword-based attack success rate (ASR) (Zou et al., 2023). This metric focuses on detecting whether the predefined keywords exist in the responses from LLMs. Such keywords typically include phrases like “I am sorry,” “As a responsible AI,” etc. Given the input pair $<J_{i},Q_{i}>$ and its corresponding response $R_{i}$, if none of the keywords from the predefined list are present in $R_{i}$, we consider that the attack has not been refused by the LLM and is successful for the corresponding sample. The keywords used for evaluations can be found in Appendix 6. The second metric is the GPT recheck attack success rate (Recheck). As we notice that sometimes the LLM does not directly refuse to answer the malicious inquiries, but delivers off-topic content. Alternatively, the LLM may reply to the malicious inquiry with added advice, such as alerting users that the request is potentially illegal or unethical. These instances might cause the ASR to be imprecise. In this context, we employ the LLM to determine if a response is essentially answering the malicious query, as demonstrated in Alg. 10. In both metrics, we report the final success rate calculated by $I_{\text{success}}/I_{\text{total}}$. For stealthiness, we use standard Sentence Perplexity (PPL) evaluated by GPT-2 as the metric.
Models. We use three open-sourced LLMs, including Vicuna-7b (Chiang et al., 2023), Guanaco-7b (Dettmers et al., 2023), and Llama2-7b-chat Touvron et al. (2023), to evaluate our method. In addition, we also use GPT-3.5-turbo (OpenAI, 2023a) to further investigate the transferability of our method to close-sourced LLMs. Additional details of our evaluations can be found in Appendix D.
@@ 4.2 Results
Effectiveness against defense. As suggested by Jain et al. (2023), we evaluate both our method and the baselines in the context against the defense method, a perplexity defense. This defense mechanism sets a threshold based on requests from the AdvBench dataset, rejecting any input message that surpasses this perplexity threshold. As demonstrated in Tab. 3, the perplexity defense significantly undermines the effectiveness of the token-level jailbreak attack, i.e., GCG attack. However, the semantically meaningful jailbreak prompts AutoDAN (and the original handcrafted DAN) is not influenced. These findings underscore the capability of our method to generate semantically meaningful content similar to benign text, verifying the stealthiness of our method.
Transferability. We further investigate the transferability of the proposed method. Following the definitions in adversarial attacks, transferability refers to in what level the prompts produced to jailbreak one LLM can successfully jailbreak another model (Papernot et al., 2016). We conduct the evaluations by taking the jailbreak prompts with their corresponding requests and targeting another LLM. The results are shown in Tab. 2. AutoDAN exhibits a much better transferability in attacking the black-box LLMs compared with the baseline. We speculate that the potential reason is the semantically meaningful jailbreak prompts may be inherently more transferable than the methods based on tokens’ gradients. As GCG-like method directly optimizes the jailbreak prompt by the gradient information, it is likely for the algorithm to get relatively overfitting in the white-box model. On the contrary, since lexical-level data such as words usually cannot be updated according to specific gradient information, optimizing at the lexical-level may make it easier to generate the more general jailbreak prompts, which may be common flaws for language models. A phenomenon that can be taken as evidence is the example shared in (Zou et al., 2023), where the authors find that using a cluster of models to generate jailbreak prompts obtains higher transferability and may produce more semantically meaningful prompts. This may support our hypothesis that the semantically meaningful jailbreak prompts are usually more transferable inherently (but meanwhile more difficult to optimize).
@@
Ablation Studies We evaluate the importance of our proposed modules in AutoDAN including the (1) DAN initialization (Sec. 3.3), (2) LLM-based Mutation (Sec. 3.5.1), and (3) the design of Hierarchical GA (Sec. 3.5). For AutoDAN-GA without DAN Initialization, we employ a prompt of a comparable length, instructing the LLM to behave as an assistant that responds to all user queries.
@@
The results are presented in Tab. 5. These results show that the modules we introduced consistently enhance the performance compared to the vanilla method. The efficacy observed with AutoDAN-GA substantiates our approach of employing genetic algorithms to formulate jailbreak prompts, validating our initial “Automatic” premise. The DAN Initialization also results in considerable improvements in both attack performance and computational speed. This is attributed to the provision of an appropriate initial space for the algorithm to navigate. Moreover, if an attack is detected as a success more quickly, the algorithm can terminate its iteration earlier and reduce computational cost. The improvements realized through DAN Initialization resonate with our second premise of “Utilizing handcrafted jailbreak prompts as prototypes.” Collectively, these observations reinforce the soundness behind the proposed AutoDAN. In addition, the LLM-based mutation yields significant improvements compared to the vanilla method, which employs basic symptom replacement. We believe that the results affirm the LLM-based mutation’s ability to introduce meaningful and constructive diversity, thereby enhancing the overall optimization process of the algorithm. The final enhancements stem from the hierarchical design. Given the effectiveness demonstrated in the original design, the hierarchical approach augments the search capability of the algorithm, allowing it to better approximate the global optimum. Furthermore, we also evaluate the effectiveness of our attack against the GPT-3.5-turbo-0301 model service by OpenAI. We use the jailbreak generated by Llama2 for testing. From results shown in Tab 5, we observe that our method can successfully attack the GPT-3.5 model and achieves superior performance compared with the baseline.

### v2

@@ 1 Introduction
However, there are two limitations of existing jailbreak methods: Firstly, automatic attacks like GCG Zou et al. (2023) inevitably request a search scheme guided by the gradient information on tokens. Although it provides a way to automatically generate jailbreak prompts, this leads to an intrinsic drawback: they often generate jailbreak prompts composed of nonsensical sequences or gibberish, i.e., without any semantic meaning (Morris et al., 2020). This severe flaw makes them highly susceptible to naive defense mechanisms like perplexity-based detection. As recent studies (Jain et al., 2023; Alon & Kamfonas, 2023) have demonstrated, such straightforward defense can easily identify these nonsensical prompts and completely undermine the attack success rate of the GCG attack. Secondly, despite the fact that manual attacks can discover stealthiness jailbreak prompts, the jailbreak prompts are often handcrafted by individual LLM users, therefore facing scalability and adaptability challenges. Moreover, such methods may not adapt quickly to updated LLMs, reducing their effectiveness over time (Albert, 2023; ONeal, 2023). Hence, a natural question emerges: “Is it possible to automatically generate stealthy jailbreak attacks? ”
@@ 3.1 Preliminaries
Genetic algorithms. Genetic Algorithms (GAs) are a class of evolutionary algorithms inspired by the process of natural selection. These algorithms serve as optimization and search techniques that emulate the process of natural evolution. GA starts with an initial population of candidate solutions (namely population initialization). Based on fitness evaluation, this population evolves with specific genetic policies, such as crossover and mutation. The algorithm concludes when a termination criterion is met, which could be reaching a specified number of generations or achieving a desired fitness threshold. The GAs can be abstracted as:
Algorithm 1 Genetic Algorithm
1:Initialize population with random candidate solutions (Sec. 3.2)
2:Evaluate fitness of each individual in the population (Sec. 3.3)
3:while termination criteria not met (Sec. 3.5do
4:     Conduct genetic policies to create offspring (Sec. 3.4)
5:     Evaluate fitness of offspring (Sec. 3.3)
6:     Select individuals for the next generation
7:end while
8:return best solution found
In this section, we introduce our design on the highlighted key components, i.e., population initialization (Sec. 3.2), fitness evaluation (Sec. 3.3), genetic policies (Sec. 3.4), termination criterion (Sec. 3.5) in their corresponding subsections.
@@ 3.2 Population Initialization
Initialization policy plays a pivotal role in genetic algorithms because it can significantly influence the algorithm’s convergence speed and the quality of the final solution. To design an effective initialization policy for AutoDAN, we should bear in mind two key considerations: 1) The prototype handcrafted jailbreak prompt has already demonstrated efficacy in specific scenarios, making it a valuable foundation; thus, it is imperative not to deviate too far from it. 2) Ensuring the diversity of the initial population is crucial, as it prevents premature convergence to sub-optimal solutions and promotes a broader exploration of the solution space. To preserve the basic features of the prototype handcrafted jailbreak prompt while also promoting diversification, we employ LLMs as the agents responsible for revising the prototype prompt, as illustrated in Alg. 5. The rationale behind this scheme is that the modifications proposed by LLM can preserve the inherent logical flow and meaning of the original sentences, while simultaneously introducing diversity in word selection and sentence structure.
3.3 Fitness Evaluation
Since the goal of jailbreak attacks can be formulated as Eq. 2, we can directly use a function that calculates this likelihood for evaluating the fitness of the individuals in genetic algorithms. Here, we adopt the log-likelihood that was introduced by Zou et al. (2023) as the loss function, namely, given a specific jailbreak prompt $J_{i}$, the loss can be calculated by:
@@ 3.3 Fitness Evaluation
3.4 Genetic Policies
3.4.1 AutoDAN-GA
Based on the initialization scheme and fitness evaluation function, we can further design the genetic policies to conduct the optimization. The core of the genetic policies is to design the crossover and mutation functions. By using a basic multi-point crossover scheme as the genetic policy, we can develop our first version of genetic algorithm, i.e., AutoDAN-GA. We provide the detailed implementation of AutoDAN-GA in Appendix C since, here, we hope to discuss how to formulate more effective policies for handling the structural discrete text data, by using its intrinsic characteristics.
@@ 3.4.2 AutoDAN-HGA
Given the population that is initialized by Alg. 5, the proposed AutoDAN will first evaluate the fitness score for every individual in the population following Eq. 3. After the fitness evaluation, the next step is to choose the individuals for crossover and mutation. Let’s assume that we have a population containing $N$ prompts. Given an elitism rate $\alpha$, we first allow the top $N*\alpha$ prompts with the highest fitness scores to directly proceed to the next iteration without any modification, which ensures the fitness score is consistently dropping. Then, to determine the remaining $N-N*\alpha$ prompts needed in the next iteration, we first use a selection method that chooses the prompt based on its score. Specifically, the selection probability for a prompt $J_{i}$ is determined using the softmax:
@@
After the selection process, we will have $N-N*\alpha$ “parent prompts” ready for crossover and mutation. Then for each of these prompts, we perform a multi-point crossover at a probability $p_{crossover}$ with another parent prompt. The multi-point crossover [^1] Since the multi-point crossover is straightforward, we defer the detailed description to Appendix C. scheme can be summarized as exchanging the sentences of two prompts at multiple breakpoints. Subsequently, the prompts after crossover will be conducted a mutation at a probability $p_{mutation}$. We let the LLM-based diversification introduced in Alg. 5 to also serve as the mutation function due to its ability to preserve the global meaning and introduce diversity. We delineate the aforementioned process in Alg. 7. For the $N-N*\alpha$ selected data, this function returns $N-N*\alpha$ offsprings. Combining these offsprings with the elite samples that we preserve, we will get $N$ prompts in total for the next iteration.
@@
At the sentence level, the search space is primarily around the word choices. After scoring each prompt using the fitness score introduced in Eq. 3, we can assign the fitness score to every word present in the corresponding prompt. Since one word can appear in multiple prompts, we set the average score as the final metric to quantify the significance of each word in achieving successful attacks. To further consider the potential instability of the fitness score in the optimization process, we incorporate a momentum-based design into the word scoring, i.e., deciding the final fitness score of a word based on the average number of the score in current iteration and the last iteration. As detailed in Alg. 8, after filtering out some common words and proper nouns (line 4) , we can obtain a word score dictionary (line 6). From this dictionary, we choose the words with top $K$ scores to replace their near-synonym in other prompts, as described in Alg. 9.
3.5 Termination Criteria
To ensure the effectiveness and efficiency of AutoDAN, we adopt termination criteria combined with a max iteration test and refusal signals test. If the algorithm has exhausted the maximum number of iterations or no keyword in a set list $L_{refuse}$ is detected in the LLM’s top $K$ words of its response, AutoDAN will terminate and return the current optimal jailbreak prompt, which has the highest fitness score. Finally, the AutoDAN-HGA can be delineated in Alg. 2.
@@ 3.5 Termination Criteria
1:Input Prototype jailbreak prompt $J_{p}$, keyword list $L_{refuse}$, and hyper-parameters
2:Initialize population with LLM-based Diversification by Alg. 5
3:while model responses have word in $L_{refuse}$ and iteration not exhausted do
4:     for $iteration$ in sentence-level iterations do
@@
6:         Calculate momentum word score by Alg. 8
7:         Update sentences in each prompt by Alg. 9
8:     for $iteration$ in paragraph-level iterations do
9:         Evaluate the fitness score of each individual in population following Eq. 3
10:         Select elite and parent prompts following Eq.4
11:         Conduct crossover and mutation on parent prompts by Alg. 7
12:end while
13:return Optimal jailbreak prompt $J_{max}$ with highest fitness score
@@ 4.1 Experimental Setups
Metrics. We use two metrics to evaluate the performance of jailbreak methods. The first metric is the keyword-based attack success rate (ASR) (Zou et al., 2023). This metric focuses on detecting whether the predefined keywords exist in the responses from LLMs. Such keywords typically include phrases like “I am sorry,” “As a responsible AI,” etc. Given the input pair $<J_{i},Q_{i}>$ and its corresponding response $R_{i}$, if none of the keywords from the predefined list are present in $R_{i}$, we consider that the attack has not been refused by the LLM and is successful for the corresponding sample. The keywords used for evaluations can be found in Appendix 6. The second metric is the GPT recheck attack success rate (Recheck). As we notice that sometimes the LLM does not directly refuse to answer the malicious inquiries, but delivers off-topic content. Alternatively, the LLM may reply to the malicious inquiry with added advice, such as alerting users that the request is potentially illegal or unethical. These instances might cause the ASR to be imprecise. In this context, we employ the LLM to determine if a response is essentially answering the malicious query, as demonstrated in Alg. 11. We provide comprehensive evaluations about the Recheck metric in Appendix F. In both metrics, we report the final success rate calculated by $I_{\text{success}}/I_{\text{total}}$. For stealthiness, we use standard Sentence Perplexity (PPL) evaluated by GPT-2 as the metric.
Models. We use three open-sourced LLMs, including Vicuna-7b (Chiang et al., 2023), Guanaco-7b (Dettmers et al., 2023), and Llama2-7b-chat Touvron et al. (2023) without system prompt, to evaluate our method. We also use GPT-3.5-turbo (OpenAI, 2023a) to further investigate the transferability of our method to close-sourced LLMs. Additional details are in Appendix D.
@@ 4.2 Results
Effectiveness against defense. As suggested by Alon & Kamfonas (2023); Jain et al. (2023), we evaluate both our method and the baselines in the context against the defense method, a perplexity defense. This defense mechanism sets a threshold based on requests from the AdvBench dataset, rejecting any input message that surpasses this perplexity threshold. As demonstrated in Tab. 3, the perplexity defense significantly undermines the effectiveness of the token-level jailbreak attack, i.e., GCG attack. However, the semantically meaningful jailbreak prompts AutoDAN (and the original handcrafted DAN) is not influenced. These findings underscore the capability of our method to generate semantically meaningful content similar to benign text, verifying the stealthiness of our method. Additionally, we also evaluate our method on other defenses in Jain et al. (2023), including paraphrasing and adversarial training, and share the results in Appendix I.
Transferability. We further investigate the transferability of our method. Following the definitions in adversarial attacks, transferability refers to in what level the prompts produced to jailbreak one LLM can successfully jailbreak another model (Papernot et al., 2016). We conduct the evaluations by taking the jailbreak prompts with their corresponding requests and targeting another LLM. The results are shown in Tab. 2. AutoDAN exhibits a much better transferability in attacking the black-box LLMs compared with the baseline. We speculate that the potential reason is the semantically meaningful jailbreak prompts may be inherently more transferable than the methods based on tokens’ gradients. As GCG-like method directly optimizes the jailbreak prompt by the gradient information, it is likely for the algorithm to get relatively overfitting in the white-box model. On the contrary, since lexical-level data such as words usually cannot be updated according to specific gradient information, optimizing at the lexical-level may make it easier to generate the more general jailbreak prompts, which may be common flaws for language models. A phenomenon that can be taken as evidence is the example shared in (Zou et al., 2023), where the authors find that using a cluster of models to generate jailbreak prompts obtains higher transferability and may produce more semantically meaningful prompts. This may support our hypothesis that the semantically meaningful jailbreak prompts are usually more transferable inherently (but more difficult to optimize).
@@
Ablation Studies We evaluate the importance of our proposed modules in AutoDAN including the (1) DAN initialization (Sec. 3.2), (2) LLM-based Mutation (Sec. 3.4.2), and (3) the design of Hierarchical GA (Sec. 3.4.2). For AutoDAN-GA without DAN Initialization, we employ a prompt of a comparable length, instructing the LLM to behave as an assistant that responds to all user queries. In addition, we investigate the efficiency of LLM-based mutation scheme, and another mutation method that utilize simple synonym replacement in Appendix H.
@@
The results are presented in Tab. 5. These results show that the modules we introduced consistently enhance the performance compared to the vanilla method. The efficacy observed with AutoDAN-GA substantiates our approach of employing genetic algorithms to formulate jailbreak prompts, validating our initial “Automatic” premise. The DAN Initialization also results in considerable improvements in both attack performance and computational speed. This is attributed to the provision of an appropriate initial space for the algorithm to navigate. Moreover, if an attack is detected as a success more quickly, the algorithm can terminate its iteration earlier and reduce computational cost. The improvements realized through DAN Initialization resonate with our second premise of “Utilizing handcrafted jailbreak prompts as prototypes.” Collectively, these observations reinforce the soundness behind the proposed AutoDAN. In addition, the LLM-based mutation yields significant improvements compared to the vanilla method, which employs basic symptom replacement. We believe that the results affirm the LLM-based mutation’s ability to introduce meaningful and constructive diversity, thereby enhancing the overall optimization process of the algorithm. The final enhancements stem from the hierarchical design. Given the effectiveness demonstrated in the original design, the hierarchical approach augments the search capability of the algorithm, allowing it to better approximate the global optimum. Furthermore, we also evaluate the effectiveness of our attack against the GPT-3.5-turbo-0301 model service by OpenAI. We use the jailbreak generated by Llama2 for testing. From results shown in Tab. 5, we observe that our method can successfully attack the GPT-3.5 model and achieves superior performance compared with the baseline. We also share the attack performance on GPT-4 in Appendix G.